# Trajectory-Based Multi-Objective Hyperparameter Optimization for Model Retraining

### v1

@@ 6.1 Experiments on Synthetic Simulations
Figure 2: Box plots for 20 synthetic problems derived from ZDT1, ZDT2, DTLZ1, DTLZ2, and DTLZ7. Each algorithm runs for 20 independent trials. The logarithm of HV difference is computed at the end of each trial and is normalized between $[0,1]$.
@@
We select five widely used multi-objective benchmark problems from ZDT [47] and DTLZ [11] test suites with $d=5$ and add Gaussian noise with a standard deviation of $1\%$ of the range of each objective. Each subplot of Figure 2 depicts the box plots over four synthetic problems with different trajectory characteristics derived from the same multi-objective benchmark. For instance, “ZDT1(M-P)” refers to the synthetic test problem derived from ZDT1 with the first objective multiplied by the monotonic curve function and the second objective by the periodic curve function. It can be observed that with a maximum budget of $150$ iterations (or the number of times the iterative learning procedure is executed), TMOBO consistently achieves the lowest HV difference among all synthetic problems, and the solutions obtained by TMOBO generally dominate a large proportion of those obtained by alternative algorithms that do not exploit trajectory information. After being enhanced by trajectory observations, the performance of two HV-based methods, $q$NEHVI-T and $q$EHVI-T, is comparable to TMOBO on ZDT2(M-M) and ZDT2(M-Q) while ParEGO-T has the worst performance. These findings indicate that the trajectory information is beneficial to the optimization of EMOHPO and that TMOBO is able to maintain its advantages when processing trajectories with even complicated characteristics. See Appendix D for more experimental results.
6.2 Experiments On Hyperparameter Tuning Benchmarks
@@
The performance of TMOBO is further examined on five different Multi-Layer Perceptron (MLP) hyperparameter tuning tasks obtained from LCBench [46], each of which aims to optimize five hyperparameters (i.e., learning rate, momentum, weight decay, max dropout rate, and max number of units) by minimizing validation loss and training cost simultaneously on a specific training dataset. As the previous work did [14, 26, 10, 31], we utilize the surrogate version of LCBench implemented in YAHPO Gym [32] and HPOBench [12], where the performance metrics (i.e., objectives) are approximated by a high-quality surrogate, in order to avoid biased implementation errors, to run more tests, and to make sure that the experiments are reproducible in any environments (see Appendix C.2). On each dataset, we observe epoch-wise model performance for any feasible hyperparameter settings up to $50$ epochs.
@@
For each task, Figure 3 plots the performance of each algorithm against time over 20 independent replications. In our analysis, we exclude the algorithm computation overhead and prioritize the cumulative model training time as the latter is the most computationally expensive part and typically dominates the total time spent on hyperparameter optimization. This ensures a clearer understanding of how optimization efforts scale with the model complexity, thereby facilitating a more insightful comparison across different models and optimization methods. Besides the enhanced MOBO methods, we also implement a variant of TMOBO, named TMOBO-nES, which does not have an early stopping mechanism and trains each sampled setting for the maximum number of epochs. We note that TMOBO and TMOBO-nES spend more time in initialization as they train the initial settings thoroughly to capture the trajectory characteristics at the initial stage. Despite the initialization, given the same wall-clock time budget, TMOBO achieves significantly lower HV difference than the other enhanced MOBO methods across all tasks. In the meantime, the average performance of TMOBO surpasses its variant TMOBO-nES, which shows the advantage of using an early stopping mechanism. While TMOBO-nES aims to learn better about trajectory characteristics through full model training, the early stopping mechanism enables TMOBO to save effort by terminating non-contributed training to explore more regions of interest.
@@ 7 Conclusions
In this study, we consider multi-objective hyperparameter optimization with iterative learning procedures. Our interest centers on how trajectory information affects the distribution of trade-offs and on how to leverage this information to perform an effective and efficient search for trade-offs. To this end, we extend the conventional MOHPO problem to EMOHPO by including the training epoch as an explicit decision variable so as to reveal the trade-offs that may occur along trajectories. These frequently overlooked trade-offs play a beneficial role in decision-making for ML model deployment with retraining schemes. As there are no algorithms specially designed to handle EMOHPO, we then propose the TMOBO algorithm to first sample the hyperparameter setting with the largest trajectory-based contribution and then determine when to early stop the training with it based on the trajectory predictions of GP models.
Through numerical experiments on synthetic simulations and hyperparameter tuning benchmarks, TMOBO has demonstrated its advantage in locating better solutions by exploiting the trajectory information compared to traditional multi-objective optimization methods. Considering that the iterative processing procedure inherent in many real-world simulations or experiments, such as drug design and material engineering, shares similar characteristics with the training of ML models, it is meaningful to explore the formulation of EMOHPO problems across a variety of practical scenarios and to extend the success of TMOBO algorithm in this study. However, optimizing the TEHVI acquisition function remains challenging. Therefore, further exploitation is needed to derive an analytical form of TEHVI or to develop more efficient approximations. Another research direction involves examining how the integration of dependent objectives or learning curves into GP modeling as well as the optimization framework can further enhance the decision-making processes.
8 Acknowledgement
We would like to thank Songhao Wang and Haowei Wang for their helpful discussions about this study. Szu Hui Ng’s work is supported in part by the Ministry of Education, Singapore (Grant: R-266-000-149-114).
Supplementary Material for
Trajectory-Based Multi-Objective Hyperparameter Optimization for Model Retraining
Appendix A Proof for Theorem 1
Recall that Theorem 1 claims that the Pareto-optimal set of $X_{Trj}^{*}\times\mathbb{T}$ is equal to the Pareto-optimal set of EMOHPO, where $X_{Trj}^{*}=\{\bm{x}\mid\bm{x}\in\mathbb{X},\exists t\in\mathbb{T},\nexists(\bm{x}^{\prime},t^{\prime})\in\mathbb{X}\times\mathbb{T},\bm{f}(\bm{x}^{\prime},t^{\prime})\prec\bm{f}(\bm{x},t)\}$. We start by establishing the following lemma.
@@ Appendix A Proof for Theorem 1
As each query pair $\bm{z}=\{\bm{x},t\}$, the Pareto-optimal set $Z^{*}$ of EMOHPO defined in Definition 2 can be equivalently expressed as,
@@
$Z^{*}$$\subseteq\{(\bm{x},t^{\prime\prime})\mid\bm{x}\in\mathbb{X},\exists t\in\mathbb{T},(\bm{x},t)\in Z^{*},t^{\prime\prime}\in\mathbb{T}\}$ (11)
=$=\{\bm{x}\mid\bm{x}\in\mathbb{X},\exists t\in\mathbb{T},(\bm{x},t)\in Z^{*}\}\times\{t^{\prime\prime}\mid t^{\prime\prime}\in\mathbb{T}\}=X_{Trj}^{*}\times\mathbb{T}.$
@@
Appendix B Additional Algorithm Details
Acquisition function plays a critical part in the BO framework but maximizing it presents inherent challenges. As discussed in Section 5.1, the TEHVI acquisition function extended from EHVI has no known analytical form and is multi-modal. Furthermore, approximating TEHVI by the MC method in (6) incurs significant computational overhead, making it hard to optimize efficiently. To this end, our algorithm TMOBO adopts a candidate search to determine new hyperparameter settings by strategic sampling instead of iterative optimization of the acquisition function. The candidate search first proposed by Regis and Shoemaker [34] aims to generate a set of random settings (termed as “candidates”) around a previously visited high-quality hyperparameter setting (termed as “center”). The new hyperparameter setting is then selected from these candidates by comparing their respective acquisition function values. This approach has demonstrated its effectiveness in global optimization [34] and has been extended to multi-objective cases in [43, 42]. The specific implementation details of the candidate search in TMOBO are provided below:
@@ Appendix B Additional Algorithm Details
$\bm{x}^{c}=\operatorname*{argmax}_{\bm{x}\in X}~{}\big{[}HV\big{(}F\mid\bm{r}\big{)}-HV\big{(}F\setminus\big{\{}\bm{y}(\bm{x},t)\big{\}}_{t=0}^{t^{\prime}_{max}}\mid\bm{r}\big{)}\big{]},$ (12)
@@ C.1 Algorithm Configurations
The Python implementation for TMOBO will be made openly available upon publication. Within the candidate search strategy of TMOBO, the initial search radius $\gamma_{max}$ associated with any visited hyperparameter setting is set to $0.2$ as recommended by [34]. Moreover, to ensure a sufficiently dense neighborhood around the center, the number of candidates $q$ generated around a center is set to $100d$. Then, in the early stopping mechanism, we determine the conservative stopping epoch in a way similar to computing the lower confidence bound and maintain a fixed value of $\beta$, which controls the confidence level, at $2.0$. Finally, at the end of each iteration, TMOBO takes an active data augmentation method to minimize the prediction uncertainty of the GP model by using only a subset of trajectory observations. To maintain the GP training efficiency, we limit the size of this subset to a maximum of $10$.
@@ C.2 Problems and Benchmarks
ZDT1 Benchmark [47]:
$\min_{\bm{x}}~{}[f_{1}(\bm{x}),f_{2}(\bm{x})],$ (13)
$f_{1}(\bm{x})=x_{1},$
$f_{2}(\bm{x})=u(\bm{x})\left[1-\sqrt{x_{1}/u(\bm{x})}\right]$
@@
ZDT2 Benchmark [47]:
$\min_{\bm{x}}~{}[f_{1}(\bm{x}),f_{2}(\bm{x})],$ (14)
$f_{1}(\bm{x})=x_{1},$
@@
DTLZ1 Benchmark [11]:
$\min_{\bm{x}}~{}[f_{1}(\bm{x}),\dots,f_{k}(\bm{x})],$ (15)
@@
DTLZ2 Benchmark [11]:
$\min_{\bm{x}}~{}[f_{1}(\bm{x}),\dots,f_{k}(\bm{x})],$ (16)
@@
DTLZ7 Benchmark [11]:
$\min_{\bm{x}}~{}[f_{1}(\bm{x}),\dots,f_{k}(\bm{x})],$ (17)
@@
Figure 4: Average log Hypervolume difference against iterations for each algorithm on synthetic problems generated from ZDT1 and ZDT2. Each algorithm runs for 20 independent trials. The shaded region indicates two standard errors of the mean.

### v2

@@ 6.1 Synthetic Simulations
We select five widely used multi-objective benchmark problems from ZDT [Zitzler et al., 2000] and DTLZ [Deb et al., 2002] test suites with $d=5$ and add Gaussian noise with a standard deviation of $1\%$ of the range of each objective. Each subplot of Figure 2 depicts the box plots over four synthetic problems with different trajectory characteristics derived from the same multi-objective benchmark. For instance, “ZDT1(M-P)” refers to the synthetic test problem derived from ZDT1 with the first objective multiplied by the monotonic curve function and the second objective by the periodic curve function. It can be observed that with a maximum budget of $150$ iterations (or the number of times the iterative learning procedure is executed), TMOBO consistently achieves the lowest HV difference among all synthetic problems, and the solutions obtained by TMOBO generally dominate a large proportion of those obtained by alternative algorithms that do not exploit trajectory information. After being enhanced by trajectory observations, the performance of two HV-based methods, $q$NEHVI-T and $q$EHVI-T, is comparable to TMOBO on ZDT2(M-M) and ZDT2(M-Q) while ParEGO-T has the worst performance. These findings indicate that the trajectory information is beneficial to the optimization of EMOHPO and that TMOBO is able to maintain its advantages when processing trajectories with even complicated characteristics.
6.2 Hyperparameter Tuning Benchmarks
@@
Five different Multi-Layer Perceptron (MLP) hyperparameter tuning tasks obtained from LCBench [Zimmer et al., 2021] are first utilized to examine the performance of TMOBO, where each task aims to optimize five hyperparameters (i.e., learning rate, momentum, weight decay, max dropout rate, and max number of units) by minimizing validation loss and training cost simultaneously on a specific training dataset. Following previous work [Falkner et al., 2018, Martinez-Cantin, 2018, Daxberger et al., 2019, Perrone et al., 2018], we utilize the surrogate version of LCBench implemented in YAHPO Gym [Pfisterer et al., 2022] and HPOBench [Eggensperger et al., 2021], where the performance metrics (i.e., objectives) are approximated by a high-quality surrogate. This approach avoids biased implementation errors, enables extensive testing, and ensures reproducibility across any environments (see Appendix C.2). On each dataset, we observe epoch-wise model performance for any feasible hyperparameter settings up to $50$ epochs.
@@
Figure 3 compares algorithm performance over 20 independent replications, measured against cumulative model training time (excluding algorithm overhead) since model training is generally more computationally expensive and dominates hyperparameter optimization costs. This ensures a clearer understanding of how optimization efforts scale with the model complexity, thereby facilitating a more insightful comparison across different models and optimization methods. Besides the enhanced MOBO methods, we also implement a variant of TMOBO, named TMOBO-nES, which does not have an early stopping mechanism and trains each sampled setting for the maximum number of epochs. We note that TMOBO and TMOBO-nES spend more time in initialization as they train the initial settings thoroughly to capture the trajectory characteristics. Despite the initialization, given the same time budget, TMOBO achieves significantly lower HV difference than the other enhanced MOBO methods across all tasks. Meanwhile, the average performance of TMOBO surpasses TMOBO-nES, which shows the advantage of using an early stopping mechanism. While TMOBO-nES aims to learn better about trajectory characteristics through full model training, the early stopping mechanism enables TMOBO to save effort by terminating non-contributed training to explore more regions of interest.
Figure 4: Average log HV difference against wall-clock time for each algorithm on the hyperparameter tuning task of MobileNetV2 on the CIFAR-10 dataset. The shaded region indicates two standard errors of the mean.
We further apply TMOBO, TMOBO-nES, and $q$NEHVI-T to a more challenging CNN-based task: tuning MobileNetV2 [Sandler et al., 2018] on CIFAR-10 image dataset by optimizing learning rate, momentum, weight decay, batch size, and training epochs. However, unlike LCBench experiments, here we measure total wall-clock time, including CNN training and algorithm overhead. Figure 4 shows that after 35,000 seconds, both TMOBO and TMOBO-nES surpass qNEHVI-T in HV difference, with TMOBO converging faster early on due to early stopping. Though TMOBO-nES eventually matches TMOBO’s performance, the latter’s efficiency in the initial phase highlights its practical advantage for computationally expensive tasks. Given the inherent complexity of this CNN-based task, we believe these results provide strong evidence for the scalability and practical applicability of TMOBO.
@@ 7 Conclusions
In this study, we consider multi-objective hyperparameter optimization with iterative learning procedures. Our interest centers on how trajectory information affects the distribution of trade-offs and on how to leverage this information to perform an effective and efficient search for trade-offs. To this end, we extend the conventional MOHPO problem to EMOHPO by including the training epoch as an explicit decision variable so as to reveal the trade-offs that may occur along trajectories. These frequently overlooked trade-offs play a beneficial role in decision-making for addressing overfitting issue and optimizing ML model deployment with retraining schemes. As there are no algorithms specially designed to handle EMOHPO, we then propose the TMOBO algorithm to first sample the hyperparameter setting with the largest trajectory-based contribution and then determine when to early stop the training with it based on the trajectory predictions of GP models.
Through experiments on synthetic simulations and hyperparameter tuning benchmarks, TMOBO has demonstrated its advantage in locating better solutions by exploiting the trajectory information compared to traditional multi-objective optimization methods. Considering that the iterative processing procedure inherent in many real-world simulations or experiments, such as drug design and material engineering, shares similar characteristics with the training of ML models, it is meaningful to explore the formulation of EMOHPO problems across a variety of practical scenarios and to extend the success of TMOBO algorithm in this study. However, optimizing the TEHVI acquisition function remains challenging. Therefore, further exploitation is needed to derive an analytical form of TEHVI or to develop more efficient approximations.
Acknowledgements.
We gratefully acknowledge the anonymous reviewers for their valuable feedback and constructive suggestions. We would also like to thank Songhao Wang and Haowei Wang for their helpful discussions regarding this study. Szu Hui Ng’s work is supported in part by the Ministry of Education, Singapore (Grant: R-266-000-149-114).
Appendix A Proof for Lemma 1
Recall that Lemma 1 claims that the Pareto-optimal set of $X_{Trj}^{*}\times\mathbb{T}$ is equal to the Pareto-optimal set of EMOHPO, where $X_{Trj}^{*}=\{\bm{x}\mid\bm{x}\in\mathbb{X},\exists t\in\mathbb{T},\nexists(\bm{x}^{\prime},t^{\prime})\in\mathbb{X}\times\mathbb{T},\bm{f}(\bm{x}^{\prime},t^{\prime})\prec\bm{f}(\bm{x},t)\}$. We start by establishing the following lemma.
@@ Appendix A Proof for Lemma 1
As each query pair $\bm{z}=\{\bm{x},t\}$, the Pareto-optimal set $Z^{*}$ of EMOHPO, i.e., the trade-offs over trajectories, can be equivalently expressed as,
@@
$Z^{*}$$\subseteq\{(\bm{x},t^{\prime\prime})\mid\bm{x}\in\mathbb{X},\exists t\in\mathbb{T},(\bm{x},t)\in Z^{*},t^{\prime\prime}\in\mathbb{T}\}$
=$=\{\bm{x}\mid\bm{x}\in\mathbb{X},\exists t\in\mathbb{T},(\bm{x},t)\in Z^{*}\}\times\{t^{\prime\prime}\mid t^{\prime\prime}\in\mathbb{T}\}=X_{Trj}^{*}\times\mathbb{T}.$ (7)
@@
Appendix B Algorithm Details
B.1 Gaussian Process Prediction in TMOBO
We begin by visualizing GP’s capability of predicting learning curve at a single iteration while running TMOBO on a real hyperparameter tuning benchmark kc1. The prediction results are illustrated in Figure B.1. Initially, even though the learning curve (depicted by orange stars) follows an uncommon shape and remains entirely unobserved, the GP model, which has been trained on data from the first 35 iterations, provides a high-quality approximation of the learning curve (depicted by the green curve) of the selected hyperparameter setting $\bm{x}$. This allows TMOBO to establish a solid foundation for trajectory prediction by leveraging multiple GP models simultaneously and computing a reliable TEHVI acquisition value before the algorithm decides to start the iterative learning process at $\bm{x}$.
As the iterative learning process progresses for the selected setting $\bm{x}$, true but noisy observations are revealed to the GP model epoch by epoch. These observations continuously enhance the predictive quality of the GP model. As shown in Figure B.1, after incorporating the revealed observations (depicted by red stars) from the learning curve, the updated GP predictions align more closely with the true data. This refinement leads to a substantial reduction in prediction uncertainty, enabling TMOBO to make more informed decisions for early stopping.
Figure B.1: An illustrative example of GP prediction obtained after running TMOBO on the kc1 hyperparameter tuning task for 35 iterations. [Left] GP prediction (green curve) of validation loss for the selected hyperparameter setting over epochs 1 to 50 before any true observations (orange stars) from its learning curve are known. [Right] Updated GP predictions after some observations (red stars) on the learning curve are revealed and incorporated into the GP model.
B.2 Candidate Search Strategy
Acquisition function plays a critical part in the BO framework but maximizing it presents inherent challenges. As discussed in Section 5.1, the TEHVI acquisition function extended from EHVI has no known analytical form and is multi-modal. Furthermore, approximating TEHVI by the MC method incurs significant computational overhead, making it hard to optimize efficiently. To this end, our algorithm TMOBO adopts a candidate search to determine new hyperparameter settings by strategic sampling instead of iterative optimization of the acquisition function. The candidate search first proposed by Regis and Shoemaker [2007] aims to generate a set of random settings (termed as “candidates”) around a previously visited high-quality hyperparameter setting (termed as “center”). The new hyperparameter setting is then selected from these candidates by comparing their respective acquisition function values. This approach has demonstrated its effectiveness in global optimization [Regis and Shoemaker, 2007] and has been extended to multi-objective cases in [Wang et al., 2023, Wang and Shoemaker, 2023]. The specific implementation details of the candidate search in TMOBO are provided below:
@@ B.2 Candidate Search Strategy
$\bm{x}^{c}=\operatorname*{argmax}_{\bm{x}\in X}\leavevmode\nobreak\ \big{[}HV\big{(}F\mid\bm{r}\big{)}-HV\big{(}F\setminus\big{\{}\bm{y}(\bm{x},t)\big{\}}_{t=0}^{t^{\prime}_{max}}\mid\bm{r}\big{)}\big{]},$ (8)
@@ C.1 Algorithm Configurations
As discussed in Section 4, the kernel of GP models in TMOBO is implemented by the product of a Matérn kernel (with smoothness parameter $\nu=2.5$) over the hyperparameter space and a temporal kernel over epochs. In real-world benchmarks, where some prior knowledge of the objective’s characteristics is available, we use informed temporal kernels: an exponential decay kernel when the objective is validation loss and a linear kernel when the objective is training cost. However, in synthetic simulations, which are designed to reflect more general settings, we default to a Matérn temporal kernel with $\nu=2.5$, providing a general-purpose modeling choice. The GP models are implemented using the GPyTorch library [Gardner et al., 2018] and fitted via standard maximum likelihood estimation with gradient-based optimization of kernel hyperparameters.
Furthermore, within the candidate search strategy of TMOBO, the initial search radius $\gamma_{max}$ associated with any visited hyperparameter setting is set to $0.2$ as recommended by [Regis and Shoemaker, 2007]. Moreover, to ensure a sufficiently dense neighborhood around the center, the number of candidates $q$ generated around a center is set to $100d$. Then, in the early stopping mechanism, we determine the conservative stopping epoch in a way similar to computing the lower confidence bound and maintain a fixed value of $\beta$, which controls the confidence level, at $2.0$. Finally, at the end of each iteration, TMOBO takes an active data augmentation method to minimize the prediction uncertainty of the GP model by using only a subset of trajectory observations. To maintain the GP training efficiency, we limit the size of this subset to a maximum of $10$.
@@ C.2 Problems and Benchmarks
ZDT1 Benchmark [Zitzler et al., 2000]:
$\min_{\bm{x}}\leavevmode\nobreak\ \leavevmode\nobreak$ $[f_{1}(\bm{x}),f_{2}(\bm{x})],$
$f_{1}(\bm{x})=x_{1},$ (9)
$f_{2}(\bm{x})=u(\bm{x})\left[1-\sqrt{x_{1}/u(\bm{x})}\right],$
@@
ZDT2 Benchmark [Zitzler et al., 2000]:
$\min_{\bm{x}}\leavevmode\nobreak\ \leavevmode\nobreak$ $[f_{1}(\bm{x}),f_{2}(\bm{x})],$
$f_{1}(\bm{x})=x_{1},$ (10)
@@
DTLZ1 Benchmark [Deb et al., 2002]:
$\min_{\bm{x}}\leavevmode\nobreak\ \leavevmode\nobreak$ $[f_{1}(\bm{x}),\dots,f_{k}(\bm{x})],$
@@
DTLZ2 Benchmark [Deb et al., 2002]:
$\min_{\bm{x}}\leavevmode\nobreak\ \leavevmode\nobreak$ $[f_{1}(\bm{x}),\dots,f_{k}(\bm{x})],$
@@
DTLZ7 Benchmark [Deb et al., 2002]:
$\min_{\bm{x}}\leavevmode\nobreak\ \leavevmode\nobreak$ $[f_{1}(\bm{x}),\dots,f_{k}(\bm{x})],$